OpenStudy (anonymous):

Find distance between given points (-4sqrt(3), 3sqrt(5)) and (-sqrt(3), sqrt(5)) I don't know how to work with the square roots here PLEASE HELP

OpenStudy (aum):

Distance = \(\sqrt{x_2-x_1)^2 + (y_2-y_1)^2}\) \[ x_2 - x_1 = -\sqrt{3} - (-4\sqrt{3}) = -\sqrt{3} +4\sqrt{3} = \sqrt{3}(-1+4) = 3\sqrt{3} \\ y_2 - y_1 = \sqrt{5} - 3\sqrt{5} = \sqrt{5}(1-3) = -2\sqrt{5} \\ \text{Distance = } \sqrt{x_2-x_1)^2 + (y_2-y_1)^2} = \sqrt{(3\sqrt{3})^2 + (-2\sqrt{5})^2} = \\ \sqrt{27 + 20} = \sqrt{47}\\ \]

OpenStudy (anonymous):

what did you do to get sqrt(27+20)?

OpenStudy (anonymous):

oh i see nvm